When a system of acetone undergoes a phase change from liquid to vapor, the entropy change (ΔS) for the system is 95 J/K and it requires 31.3 kJ of energy to be transferred into the system from the surroundings.

(CH3)2CO (l) → (CH3)2CO (g) ΔH = 31.3kJ

a. Calculate the overall entropy change (ΔSuniverse) for this process at 65.0°C.

b. Briefly discuss whether or not this process would be spontaneous or nonspontaneous at this temperature.
Answer #1

We know that 1 Stotal = D Spysen + DS surrounding ] A Stotal is the entropy change of Gniverse only for is a reaction et cons(W Given. Srystin = 95J/k = 95 X103 kJ/K. Dh= 31.3 kJ T= 65°C = 65+ 273.15 = 338.15 K. Sunivers = DSpor = DS nystem - AH hersWe derived A total = D D > TDS total = TDS-DH Systen we know DG DH-TDS Thus, ITDS totul = - ] uince DG is positive as we havei have also derived the desired formula used .

If you have any query regarding this you may ask in comments.